OpenStudy (anonymous):

what anual rate of interest is required to triple an investment in 12 yrs a) 4.794% b)9.578% c)5.946% d)9.155%

OpenStudy (anonymous):

compounded how?

OpenStudy (anonymous):

i think A=P(1+r/n)^ nt

OpenStudy (anonymous):

are u sure b) is 9.578 and not 9.587?

OpenStudy (anonymous):

compounded anually, semianually, monthly, daily, contiously?

OpenStudy (anonymous):

contiously?

OpenStudy (anonymous):

linhphan is b) 9.587 or 9.578

OpenStudy (anonymous):

Okay, let's solve it that way= \[A=Pe^{rt}\] since we want triple \[3P=Pe^{rt}\] \[3=e^{rt}\] now plug in number \[3=e^{r12}\]

OpenStudy (anonymous):

ln 3=12r ln e ln 3= 12 r ln 3 /12 =r

OpenStudy (anonymous):

this is what I meant \[\frac{ln 3}{12}\]

OpenStudy (anonymous):

that gives 9.155%

OpenStudy (anonymous):

So, in this kind of problem, works is much more important than mere answer